# CREStE: Scalable Mapless Navigation with Internet Scale Priors and Counterfactual Guidance

### v1

@@ VI-B Training Dataset
We collect a robot dataset with 3 hours of expert navigation demonstrations spread across urban parks, downtown centers, residential neighborhoods, and college campuses. Our dataset consists of synchronized image LiDAR observation pairs and ground truth robot poses computed using LeGO-LOAM [42], a LiDAR-based SLAM algorithm. We train all methods on the same dataset for 150 epochs or until convergence.
@@ VI-D Baselines
We test six variations of CREStE to evaluate the effect of different observation encoders and sensor modalities. These consist of 1) CREStE-RGB, the same RGB encoder but using only monocular RGB images, 2) CREStE-STEREO, the same RGB encoder with a stereo processing backbone that fuses features from stereo RGB images, 3) CREStE-RGB-Frozen, the same depth prediction head but with a frozen Dinov2 [28] encoder that only uses monocular RGB images. The remaining variations ablate our methodological contributions: 4) CREStE-cfs, our model trained without counterfactual demonstrations, but otherwise identical to the original, 5) CREStE-st, our model trained without the BEV inpainting backbone, but otherwise identical to the original, 6) CREStE-cfs-st, our model trained without counterfactual demonstrations or the BEV inpainting backbone. For architectures 5 and 6, we directly pass the unstructured BEV feature map $z_{\text{bev, splat}}$ to the reward function $r_{\phi}$ and train the splat module $f_{\text{splat}}$ using expert demonstrations. We still freeze the RGB-D backbone $f_{\text{rgbd}}$ when training $r_{\phi}$. For additional details regarding the CREStE architecture variations, we refer readers to Appendix Sec. X-C2. Next, we will describe high-level implementation details for each non-CREStE baseline.
@@ VI-E Quantitative Results and Analysis
Comparing CREStE against all other baselines in Table I, we find that our approach achieves superior performance in all metrics for both seen and unseen environments. We present qualitative analysis for each learned baseline in Appendix Sec. X-C. Quantitatively, PACER+G, the next best approach, requires two times more interventions than CREStE in seen environments and five times more interventions in unseen environments. Furthermore, we find that PIVOT, our VLM-based navigation baseline, performs poorly relative to other methods, corroborating our claim that VLMs and LLMs are not well attuned for urban navigation despite containing internet-scale priors. We hypothesize this is because navigation requires identifying which priors are most important for navigation, a task seen rarely by VLMs and LLMs during pre-training. Notably, we achieve an intervention-free traversal in one unseen environment (Hemphill Park), a residential park with diverse terrains, narrow curb gaps, and crosswalk markings. From these findings, we conclude that CREStE is remarkably more generalizable for mapless urban navigation than existing approaches.
In Table II, our approach performs most favorably using RGB and LiDAR inputs, followed by stereo RGB and monocular RGB respectively. This occurs because LiDAR offers robustness to severe lighting artifacts like lens flares that affect costmap prediction quality. We present qualitative examples of this in Fig. 10. Interestingly, we find that distilled Dinov2 features perform favorably compared to frozen Dinov2 features. Prior works corroborate this observation [51, 52] and verify that Dinov2 features contain positional embedding artifacts that disrupt multi-view consistency. We postulate that our distillation objective enables our RGB-D encoder $f_{\text{rgbd}}$ to focus on learning artifact-free features while the semantic decoder $f_{\text{semantic}}$ learns how to reconstruct these artifacts. For additional analysis comparing these architectural variations, we refer readers to Appendix Sec. X-C1.
@@
We compare CREStE against CREStE-cfs, the exact same approach but without using counterfactuals to train the reward function. On average, we find in Table I that using counterfactuals reduces the number of interventions by 70% in seen environments and 69% in unseen environments. Both of these baselines distill the same features for navigation from VFMs, thus vindicating our claim that even with sufficiently informative perceptual representations, it is important to leverage our counterfactual-based objective to properly identify and reason about the most salient features for navigation.
@@ VII Limitations and Future Work
CREStE infers reward maps using observations from a single timestep, discarding past information. This limits our approach in scenes that require temporal reasoning, such as recovering from dead ends or negotiating paths with dynamic entities like pedestrians. Extending CREStE to consider multiple observations is a promising direction that addresses these issues. Extending CREStE’s counterfactual-based IRL objective to reason about temporal dynamics is another promising direction for improving performance in crowded dynamic environments. Additionally, CREStE must be manually tuned to balance goal-reaching and reward-map aligned behavior, which poses an issue for urban environments with sporadic global guidance. Extending CREStE to jointly reason about task-specific goals along with navigation affordances will further improve real-world robustness for mapless urban navigation. While the IRL objective presented in the paper only considers expert and suboptimal demonstrations, this framework can be extended to incorporate preferences following the same derivation from the ranking perspective. This would not only allow the reward to reason about what paths are best but allow is to learn how to discriminate between two suboptimal paths should the need arise. From the perceptual aspect, a promising future direction is to extend CREStE to learn robust navigation priors from internet-scale egocentric videos.
@@ VIII Conclusion
In this paper, we introduce Counterfactuals for Reward Enhancement with Structured Embeddings (CREStE), the first framework for learning representations and policies for the full mapless navigation problem. CREStE learns robust, generalizable perceptual representations with internet-scale semantic, geometric, and entity priors by distilling features from multiple visual foundation models. We demonstrate that our perceptual representation encodes a sufficient set of factors for urban navigation, and introduce a novel counterfactual-based loss and active learning framework that teaches policies to hone in on the most important factors and infer how they influence fine-grained navigation behavior. These contributions culminate to form CREStE, a local path planning module that significantly outperforms state-of-the-art alternatives on the task of long-horizon mapless urban navigation. Through kilometer-scale real-world robot experiments, we demonstrate our approach’s effectiveness, gracefully navigating an unseen 2 kilometer urban environment with only a handful of interventions.

### v2

@@ VI-B Training Dataset
We collect a robot dataset with 3 hours of expert navigation demonstrations spread across urban parks, downtown centers, residential neighborhoods, and college campuses. Additionally, we annotate 3% of our training dataset with negative counterfactual annotations to train CREStE. Our dataset consists of synchronized image LiDAR observation pairs and ground truth robot poses computed using LeGO-LOAM [38], a LiDAR-based SLAM algorithm. We train all methods on the same dataset for 150 epochs or until convergence.
@@ VI-D Baselines
We test six variations of CREStE to evaluate the effect of different observation encoders and sensor modalities. These consist of 1) CREStE-RGB, the same RGB encoder but using only monocular RGB images, 2) CREStE-STEREO, the same RGB encoder with a stereo processing backbone that fuses features from stereo RGB images, 3) CREStE-RGB-Frozen, the same depth prediction head but with a frozen Dinov2 [25] encoder that only uses monocular RGB images. The remaining variations ablate our methodological contributions: 4) CREStE-cfs, our model trained without counterfactual demonstrations, but otherwise identical to the original, 5) CREStE-st, our model trained without the BEV inpainting backbone, but otherwise identical to the original, 6) CREStE-cfs-st, our model trained without counterfactual demonstrations or the BEV inpainting backbone. For architectures 5 and 6, we directly pass the unstructured BEV feature map $z_{\text{bev, splat}}$ to the reward function $r_{\phi}$ and train the splat module $f_{\text{splat}}$ using expert demonstrations. We still freeze the RGB-D backbone $f_{\text{rgbd}}$ when training $r_{\phi}$. For additional details regarding the CREStE architecture variations, we refer readers to Appendix Sec. X-D2. Next, we will describe high-level implementation details for each non-CREStE baseline.
@@ VI-E Quantitative Results and Analysis
Comparing CREStE against all other baselines in Table I, we find that our approach achieves superior performance in all metrics for both seen and unseen environments. We present qualitative analysis for each learned baseline in Appendix Sec. X-D. Quantitatively, PACER+G, the next best approach, requires two times more interventions than CREStE in seen environments and five times more interventions in unseen environments. Furthermore, we find that PIVOT, our VLM-based navigation baseline, performs poorly relative to other methods, corroborating our claim that VLMs and LLMs are not well attuned for urban navigation despite containing internet-scale priors. We hypothesize this is because navigation requires identifying which priors are most important for navigation, a task seen rarely by VLMs and LLMs during pre-training. Notably, we achieve an intervention-free traversal in one unseen environment (Hemphill Park), a residential park with diverse terrains, narrow curb gaps, and crosswalk markings. From these findings, we conclude that CREStE is remarkably more generalizable for mapless urban navigation than existing approaches.
In Table II, our approach performs most favorably using RGB and LiDAR inputs, followed by stereo RGB and monocular RGB respectively. This occurs because LiDAR offers robustness to severe lighting artifacts like lens flares that affect costmap prediction quality. We present qualitative examples of this in Fig. 10. Interestingly, we find that distilled Dinov2 features perform favorably compared to frozen Dinov2 features. Prior works corroborate this observation [48, 49] and verify that Dinov2 features contain positional embedding artifacts that disrupt multi-view consistency. We postulate that our distillation objective enables our RGB-D encoder $f_{\text{rgbd}}$ to focus on learning artifact-free features while the semantic decoder $f_{\text{semantic}}$ learns how to reconstruct these artifacts. For additional analysis comparing these architectural variations, we refer readers to Appendix Sec. X-D1.
@@
We compare CREStE against CREStE-cfs, the exact same approach but without using counterfactuals to train the reward function. On average, we find in Table I that using counterfactuals reduces the number of interventions by 70% in seen environments and 69% in unseen environments. Both of these baselines distill the same features for navigation from VFMs,
@@ VII Limitations and Future Work
While CREStE inherits some viewpoint-invariance from VFMs, it does not allow reward maps to be conditioned on embodiment-specific constraints, such as the traversability differences between quadrupeds and wheeled robots. An interesting future direction is applying existing research on cross-embodiment conditioning [35] to our architecture. Furthermore, CREStE infers reward maps using observations from a single timestep, limiting multi-step horizon reasoning tasks, such as recovering from dead ends or negotiating paths in constricted environments. Extending CREStE with memory and counterfactual IRL for reasoning about object dynamics are promising directions for addressing these issues. In addition, reward learning with counterfactuals is a potential bottleneck, as it requires humans for counterfactual labeling. Automated counterfactual generation with VLMs/LLMs is another promising direction to further improve scalability. Lastly, our counterfactual IRL objective can further be extended to incorporate preferences following the same derivation from the ranking perspective. This would not only allow the reward to reason about what paths are best but allow is to learn how to discriminate between two suboptimal paths should the need arise.
@@ VIII Conclusion
In this paper, we introduce Counterfactuals for Reward Enhancement with Structured Embeddings (CREStE), a scalable framework for learning representations and policies that address the open-set generalization and robustness challenges of open-world mapless navigation. CREStE learns robust, generalizable perceptual representations with internet-scale semantic, geometric, and entity priors by distilling features from multiple visual foundation models. We demonstrate that our perceptual representation encodes a sufficient set of factors for urban navigation, and introduce a novel counterfactual-based loss and active learning framework that teaches policies to hone in on the most important factors and infer how they influence fine-grained navigation behavior. These contributions culminate to form CREStE, a local path planning module that significantly outperforms state-of-the-art alternatives on the task of long-horizon mapless urban navigation. Through kilometer-scale real-world robot experiments, we demonstrate our approach’s effectiveness, gracefully navigating an unseen 2 kilometer urban environment with only a handful of interventions.